Please show me how to solve this quadratic equation step by step 2x^2+5x-1=0. Any help will be greatly appreciated.
Thanks

Answer by dkspoet(13)

2x^2+5x-1=0
x^2+(5/2)x-1/2=0
x^2+2(5/4)x+(5/4)^2-(5/4)^2-1/2=0
(x+5/4)^2-25/16-8/16=0
(x+5/4)^2=33/16
x+5/4=± sqrt(33)/4
x=-5/4 ± sqrt(33)/4